What is the slope of 2x+3y=1
Found 3 solutions by richwmiller, Greenfinch, dabanfield:
Answer by richwmiller(17219)

The slope form is y=mx+b
3y=-2x+1
y=-2/3*x+1/3
slope is m=-2/3

Answer by Greenfinch(383)

3y = 1 - 2x
y = 1/3 -( 2/3) x
so gradient is -2/3 and the y intercept is 1/3

Answer by dabanfield(803)

What is the slope of 2x+3y=1
Rewrite the equation in slope-intercept form:
y = m*x + b where m is the slope of the line.
2x + 3y = 1
3y = -2x + 1
y = (-2/3)x + 1/3
m = -2/3